A 40hz vibration generate a wave in a ripple tank if the speed of the wave in two different length of the tank are 0.8m/s and 0.5m/s calculate the different between the lengrh.
1
Expert's answer
2020-08-24T12:28:38-0400

"\lambda=\frac{v}{f}" therefore, The first distance "\lambda" 1=".\frac{0.8m/s}{40hz}=0.02m" The second distance "\lambda" 2="\frac{0.5m/s}{40}=0.013m" . The difference is "\lambda"1-"\lambda" 2 =0.02m-0.013m= .0007m≈ 0.01m.